OpenStudy (anonymous):

Medal! One Rectangle has a length of 13 1/8 inches and a perimeter of 42 1/2 inches. Write an equation and the answer and explain please. More info: The perimeter of it is equal to twice the sum of its length and width

undeadknight26 (undeadknight26):

im assuming you are trying to find the width?

undeadknight26 (undeadknight26):

Isnt it just the length subtract the perimeter?

undeadknight26 (undeadknight26):

you would do the opposite of what the original equation would do...in this case i think its adds so you would subtract your 2 known numbers to find your answers?

undeadknight26 (undeadknight26):

If it is not this please tell me maybe i misunderstood the question.

undeadknight26 (undeadknight26):

hmm...

undeadknight26 (undeadknight26):

P = 2L + 2W

undeadknight26 (undeadknight26):

that is the original equation to find the perimeter of a rectangle we have to mod it in this case.

undeadknight26 (undeadknight26):

Sorry i have to go...@ranga please help.

OpenStudy (emilylauren):

I can help :)

OpenStudy (emilylauren):

2(width)+2(length)=perimeter we know... perimeter=42.5 length=13 and 1/8 so 2(w)+2(13.125)=42.5 which simplifies to 2(w)=16 and 1/4 so w=8 and 1/8

OpenStudy (emilylauren):

let me know if you have questions or need more information :)